Glass Houses Pty has found that the cost of fitting new sliding doors is normally distributed, with an average cost per door of R12 750, and a standard deviation of R1 095. A sample of 105 new windows is randomly selected. Assuming normal distribution: What is the probability that the average cost of fitting the doors in

the sample is less than R13 000? Interpret your answer.


1
Expert's answer
2021-10-25T03:35:49-0400

"\bar{x} = 12750 \\

\sigma= 1095 \\

n=105 \\

P(\bar{x} <13000) = P(Z < \frac{13000-12750}{1095 / \sqrt{105}}) \\

= P(Z< 2.34) \\

= 0.99036"

The probability that the average cost of fitting the doors in the sample is less than R13000 is 0.99036.